A bowl contains 18 white, 10 black, and 4 red buttons, if two buttons are drawn from the bowl, what is the probability that the first one is red and the second black if the buttons are drawn without replacement
Answer by math_tutor2020(3817)


A = probability of selecting a red button
A = (number of red)/(number total)
A = (4)/(18+10+4)
A = 4/32
A = 1/8

B = probability of selecting a black button, given that event A has occurred (no replacement)
B = (number of black)/(number left after event A happens)
B = (10)/(32-1)
B = 10/31

A*B = probability of getting red, then black (no replacement)
A*B = (1/8)*(10/31)
A*B = (1*10)/(8*31)
A*B = 10/248
A*B = 5/124

Answer: 5/124